OpenStudy (anonymous):

4.A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (phi):

the | | means absolute value if the "stuff" inside is positive, then you can drop them (they are not necessary) if you have | x | and x is negative, then this is the same as -x (because -x will be positive) in general, you will get two equations for your problem |x − 0.525| = 0.01 becomes x − 0.525 = 0.01 (where x-0.525 is positive) or -(x − 0.525) = 0.01 solve both equations to find the max and min values of x

OpenStudy (anonymous):

I'm sorry but I have no idea what I'm foing do I add 0.525 to both sides @phi

OpenStudy (phi):

first, pick one of the equations let's do x − 0.525 = 0.01 yes, add +0.525 to both sides x − 0.525+0.525 = 0.01+0.525 can you simplify both sides?

OpenStudy (anonymous):

x=0.535 @phi

OpenStudy (anonymous):

| x - 0.525| = 0.01 x - 0.525 = 0.01 x = 0.01 + 0.525 x = 0.535 -| x - 0.525| = 0.01 -x + 0.525 = 0.01 -x = 0.01 - 0.525 -x = 0.515 x = - 0.515 @phi

OpenStudy (phi):

the first one is correct the second is ok up to -x = 0.010 - 0.525 one way to do this is multiply both sides and *all* terms by -1 to get x= -0.010 + 0.525 which is easier to work

OpenStudy (anonymous):

minimum is : -0.515 diameter maximum is : 0.535 diameter Am I right @phi

OpenStudy (phi):

yes for the max. redo the min